OpenStudy (anonymous):

The ordered pair (1, -5) is a solution for the system x + 2y = -9 and x - y = 6 ~ True ~ False

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Hello! You have multiple ways of doing this problem. One way is to find the actual colution of the system of (2) linear equations. The other is to determine whether or not (1,-5) satisfies both of the 2 equations. Which would you prefer?

OpenStudy (anonymous):

I would prefer to determine whether or not (1,-5) satisfies both of the 2 equations.

OpenStudy (mathmale):

Fine. Let's do that. Take the first equation, x + 2y = -9. substitute x=- and y=-5 into that equation. Is the eq'n true or false?

OpenStudy (anonymous):

So, it's now x=-2 and y=-5?

OpenStudy (mathmale):

actually, MM, we were to determine wehther or not that first equation, x+2y=-9, is true when we subst. x=1 and y=-5. If I do that, I get 1+2(-5)=-9. Do you regard this as true or false?

OpenStudy (mathmale):